Exercise 3-9 Applying Overhead; T-accounts; Journal Entries [L03-1, LO3-2, LO3-4] Harwood Company uses a job-order costing system that applies overhead cost to jobs on the basis of machine-hours. The companys predetermined overhead rate of $2.60 per machine-hour was based on a cost formula that estimates $257,400 of total manufacturing overhead for an estimated activity level of 99,000 machine-hours. Required 1. Assume that during the year the company works only 94,000 machine-hours and incurs the following costs in the Manufacturing Overhead and Work in Process accounts: Compute the amount of overhead cost that would be applied to Work in Process for the year and make the entry in your T-accounts. 2A. Compute the amount of underapplied or overapplied overhead for the year and show the balance in your Manufacturing Overhead T-account. 2B. Prepare a journal entry to close the companys underapplied or overapplied overhead to Cost of Goods Solo Complete this question by entering your answers in the tabs below Req 1 Req 2A Req 2B Compute the amount of overhead cost that would be applied to Work in Process for the year and make the entry in your T-accounts Manufacturing Overhead 40,000 9,900 79,000 46,000 8,900 64,000 Work in Process (Maintenance) (Indirect materials) (Indirect labor) (Utilities) (Insurance) (Depreciation) Balance (Direct materials) (Direct labor) (Overhead) (a) 900,000 109,000 Req 1 Req 2A>Req 1 Req 2A Req 2B Compute the amount of underapplied or overapplied overhead for the year and show the balance in your Manufacturing Overhead T-account from Req 1. manufacturing overheadReq 1 Req 2A Req 2B Prepare a journal entry to close the companys underapplied or overapplied overhead to Cost of Goods Sold. (If no entry is required for a transaction/event, select No journal entry required in the first account field. View transaction list Journal entry worksheet Record the entry to close the balance of the manufacturing overhead account to COGS account. Note: Enter debits before credits. Event General Journal Debit Credit Record entry Clear entry View general journal
Answer #1

1) T account

Manufacturing overhead Work in process
Maintenance 40000 94000*2.60 = $244400 Work in process Direct material 900000      
Indirect material 9900 Direct labor 109000
Indirect labor 79000 Manufacturing overhead 244400
Utilities 46000
Insurance 8900
Depreciation 64000
Balance 3400

2) Under applied manufacturing overhead = (247800-244400) = $3400

3) Journal entry

Date account and explanation debit credit
Cost of goods sold 3400
Manufacturing overhead 3400
(To record under applied)